To what temperature must a given mass of nitrogen at 0oC heated so that both its volume and pressure will be doubled?

1
Expert's answer
2022-05-05T02:39:02-0400

PV=nRT

"T_1=0\u00b0C=273 K"

"P_2=2P_1"

"V_2=2V_1"

"n=\frac{P_1V_1}{RT_1}"

"T_2=\frac{P_2V_2}{nR}=\frac{P_2V_2RT_1}{RP_1V_1}=\frac{2P_1(2V_1)273}{P_1V_1}=1092 K=819°C"